OpenStudy (anonymous):

did I get my answers right? An expert shot hits the target 95% of the time. What are the probabilities that he will miss the target: a.for the first time on the fifteenth shot? b.at least one time during fifteen shots? c.for the second time on the eighteenth shot? a) P(X=0;14, 1/20) * P(X=1;1, 1/19) ≈ 0.3593 * 0.05 ≈ 0.0180 b) P(X=1;15, 1/20) ≈ 0.3658 c) P(X=1;1, 1/20) * P(X=1;7, 1/20) ≈ 0.05 * 0.2708 ≈ 0.0349

OpenStudy (vishweshshrimali5):

How did you obtain 0.3593 in a)?

OpenStudy (anonymous):

susing the formula for binomial distributions[ n!/(x!(n-x)!) ]* p^x *q^(n-x)

OpenStudy (vishweshshrimali5):

Okay

OpenStudy (vishweshshrimali5):

But, in P(X=1;1, 1/19) what does 1/19 mean ?

OpenStudy (vishweshshrimali5):

I think p represents probability of success which is 95/100 i.e. 19/20 and q represents probability of failure which is 5/100 i.e. 1/20 . Is this right ?

OpenStudy (anonymous):

yes that's right

OpenStudy (anonymous):

and x is the # of 'successes'

OpenStudy (vishweshshrimali5):

You should then have used directly this for (a) required answer is of getting P(X = 14) in 14 trials and then multiplying with P(X=0) i.e. no success in 1 trial (which is the 15th one).

OpenStudy (vishweshshrimali5):

Here by P(X=k) I mean probability of getting exactly k successes in n trials.

OpenStudy (vishweshshrimali5):

So, using my method of representation: (A) P(X=14;14)*P(X=0;1)

OpenStudy (anonymous):

I was Using P(X=0; 14, 1/20) because I thought that the 'success' here was not hitting a target

OpenStudy (vishweshshrimali5):

So for (A) the answer should be: \[\large{\left(\begin{matrix}14 \\14\end{matrix}\right)*\left( \cfrac{ 19 }{ 20 } \right)^{14}*\left( \cfrac{ 1 }{ 20 } \right)^{0}}\]

OpenStudy (anonymous):

or is that just semantics?

OpenStudy (vishweshshrimali5):

Well you can take success for that also, but your main mistake was in this \(\color{blue}{\text{Originally Posted by}}\) @pblaesre a) P(X=0;14, 1/20) * \(\large{P(X=1;1, 1/19)}\) ≈ 0.3593 * 0.05 ≈ 0.0180 \(\color{blue}{\text{End of Quote}}\)

OpenStudy (vishweshshrimali5):

In the enlarged part, the probability for failure should have been 19/20

OpenStudy (anonymous):

yes I see that thanks,

OpenStudy (vishweshshrimali5):

For (B) it should be : Req. probability P = 1 - P(X=0;15) Remember here I have used your notations.

OpenStudy (vishweshshrimali5):

It means the required probability should be 1 - probability of not missing any target i 15 trials.

OpenStudy (anonymous):

shouldn't it be: P = 1 - P(X=1;15) since it is "b.at least one time during fifteen shots?"

OpenStudy (vishweshshrimali5):

At least one's conjugate is not even one

OpenStudy (vishweshshrimali5):

So, it should be P = 1 - P(X=0;15)

OpenStudy (vishweshshrimali5):

Did you get till this step ?

OpenStudy (anonymous):

Sorry I'm still mentally working on the first one, I want to make sure I have it right

OpenStudy (vishweshshrimali5):

Okay

OpenStudy (anonymous):

I'm a bit slow haha

OpenStudy (vishweshshrimali5):

Its okay

OpenStudy (vishweshshrimali5):

For missing first time on 15th shot you have to ensure that no shot is missed in first 14 trials and then the shot must miss on 15th trial. Right ?

OpenStudy (anonymous):

I think I got that

OpenStudy (anonymous):

(A) [P(X=14;14, 19/20)*P(X=0;1, 19/20)] = [14P14 (19/20)^(14)(1/20)^(0) * 0P1 (19/20)^(0)(1/20)^(1)] ≈ 0.0244

OpenStudy (vishweshshrimali5):

Its actually 14C14 and 1C0

OpenStudy (anonymous):

so the first 14 are successes and the last on is a failure

OpenStudy (vishweshshrimali5):

Yes if you mean by success not missing and by failure - missing

OpenStudy (anonymous):

yes

OpenStudy (vishweshshrimali5):

Good.

OpenStudy (anonymous):

the book i'm reading says it's a type of permutaion ( a permutation of similar objects) butit looks like a combination

OpenStudy (vishweshshrimali5):

SO your answer for (A) would be : \[\large{\left(\begin{matrix}14 \\ 14\end{matrix}\right)*(\cfrac{19}{20})^{14} * \left(\begin{matrix}1 \\ 0\end{matrix}\right)*(\cfrac{1}{20})^{1}}\]

OpenStudy (vishweshshrimali5):

Yes it is a combination

OpenStudy (vishweshshrimali5):

Here, \[\left(\begin{matrix}n \\ r\end{matrix}\right) = \cfrac{n!}{(r!)(n-r)!}\]

OpenStudy (anonymous):

yup that's what I got

OpenStudy (vishweshshrimali5):

Good

OpenStudy (vishweshshrimali5):

Then move on to (B).

OpenStudy (anonymous):

\[\left(\begin{matrix}15 \\ 0\end{matrix}\right) \left( \frac{ 19 }{ 20 } \right)^{15}\]

OpenStudy (anonymous):

is that even close?

OpenStudy (anonymous):

1-that =0.5367

OpenStudy (vishweshshrimali5):

Great work @pblaesre You really grabbed the concept.

OpenStudy (anonymous):

thanks :)

OpenStudy (vishweshshrimali5):

Lets move on to (C).

OpenStudy (vishweshshrimali5):

Shall we ?

OpenStudy (anonymous):

Is it okay if I try and check with you if i'm right?

OpenStudy (vishweshshrimali5):

That's fine by me.

OpenStudy (vishweshshrimali5):

GO ahead give it a try

OpenStudy (anonymous):

\[\left(\begin{matrix}17 \\ 16\end{matrix}\right)\left(\begin{matrix}\frac{ 19 }{ 20 } \\ \end{matrix}\right)^{16}\left(\begin{matrix}\frac{ 1 }{ 20 } \\ \end{matrix}\right)^{1} \times \left(\begin{matrix}1 \\0\end{matrix}\right)\left(\begin{matrix}\frac{ 19 }{ 20 } \\ \end{matrix}\right)^{0}\left(\begin{matrix}\frac{ 1 }{ 20 } \\ \end{matrix}\right)^{1}\] from P( X=16;17) * P(X=0;1)

OpenStudy (vishweshshrimali5):

Great work indeed @pblaesre , great work. You got that right, Perfect answer.

OpenStudy (anonymous):

thanks! this was really confusing me at first.

OpenStudy (vishweshshrimali5):

No problem. But, since I can't give more than one medal you will have to get satisfied with my words of appreciation only. Seriously, you have done great work. Keep this hard work and you will master probability in no time. Best of luck Vishwesh